What happens to the moment of inertia of a cylinder if we half its mass and double its radius? Enter a number that expresses the new inertia in terms of the old
Answer #1

Assuming your question is about solid cylinder

The Moment of Inertial of cylinder with Mass M and Radius R along its axis is given by

\frac{1}{2}MR^{2}

Let initial Moment of Inertia be I0

I_0= \frac{1}{2}MR^{2}

Now Mass is reduced to half and Radius is doubled

Let new Moment of inertia be Inew

I_{new}= \frac{1}{2}\frac{M}{2}(2R)^{2}

I_{new}= MR^{2}

I_{new}= 2\times I_{0}

\frac{I_{new}}{I_{0}}= 2